# Jina-ColBERT-v2: A General Purpose Multilingual Late Interaction Retriever

### v1

@@
Jina-ColBERT-v2: A General Purpose Multilingual Late Interaction Retriever
Rohan Jha
The University of Texas at Austin
Austin, Texas, USA
rjha@utexas.edu
&Bo Wang
Jina AI
Berlin, Germany
bo.wang@jina.ai
&More Names
Jina AI
Berlin, Germany
more.names@jina.ai
  Work done while at Jina AI
Multi-vector dense models like ColBERT have demonstrated their great utility in the retrieval space. ColBERT’s late interaction scoring approximates the joint query-document attention of cross encoders, while maintaining inference efficiency closer to that of conventional dense retrieval thanks to its bi-encoder architecture and a number of recent works optimizing the indexing and search steps. In this work we propose a number of incremental improvements to the ColBERT model architecture and training pipeline, using methods shown to work in the more mature single-vector embedding model training paradigm, particularly those that apply to heterogeneous multilingual data. Our resulting model, Jina-ColBERT-v2, exhibits strong performance across a variety of English and multilingual retrieval tasks, while reducing storage requirements by as much as half compared to previous models.
@@ 1 Introduction
Neural retrieval has gained popularity in recent years following the arrival of capable pre-trained language models (PLMs) such as devlin_bert_2019; liu2019roberta; clark_electra_2020. Sparse neural retrieval systems, such as SPLADE formal_splade_2021, represent texts as weighted bags of words. In contrast, dense retrievers encode queries and documents as high-dimensional vectors, capturing relevance signals through spatial relationships extending beyond exact term matching.
Most dense retrievers encode a query or document as a single vector, commonly the result of mean- or [CLS]-pooling over the transformer’s final layer token embeddings. In contrast, recent multi-vector retrievers such as ColBERT (khattab_colbert_2020) generalize this embedding process to maintain a (usually smaller) embedding for each token, computing relevance scores as a function of the similarities of query and document tokens instead. This approach has the benefit of remaining compatible with much of the vector similarity infrastructure that makes single-vector methods efficient, but requires more space to store an embedding per token and compute at inference time to aggregate token interactions into a single score. This "late interaction" over token embeddings achieves greater in-domain performance and tends to be more robust out-of-domain than single-vector similarity. While ColBERTv2 is trained only on English MSMARCO triplets (bajaj2016msmarco) and has a monolingual BERT backbone, making it incapable of multilingual retrieval, some previous works extend ColBERT to multilingual retrieval. (louis_colbert-xm_2024) does this by using parameter extensions for each additional language, and (lawrie_neural_2023) trains solely on machine-translated English MSMARCO data to get effective multilingual performance. However, these approaches come with significant trade-offs in terms of model usability and training data diversity. Other multilingual multi-vector models like BGE-M3 (chen_bge_2024) produce extremely large token representations that limit their practical utility for first-stage retrieval.
@@
In this work, we propose Jina-ColBERT-v2111https://huggingface.co/jinaai/jina-colbert-v2 an improved recipe for training ColBERT models with the following features:
  1. 1.
    Training with diverse weakly supervised data: Our two-stage recipe includes a large-scale contrastive tuning stage followed by a smaller-scale fine-tuning stage with supervised distillation.
  2. 2.
@@
    General multilingual performance: We train with data from a variety of high- and low-resource languages, including machine-translated training data, and show that this improves even out-of-domain multilingual performance.
  3. 3.
@@
    Inference-agnostic efficiency: We introduce multiple sizes of linear projection heads, jointly trained using the non-weight tying variant of Matryoshka Representation Loss kusupati_2022_matryoshka, enabling the selection of token embedding size at inference time with minimal performance degradation. Additionally, our flash-attention optimized backbone, Jina-XLM-RoBERTa  provides further performance boosts during inference.
@@ 2.1 Single-Vector Retrieval
Single-vector encoder models have demonstrated their potential as general-purpose embedding models across a number of downstream tasks (muennighoff-etal-2023-mteb). When used in a bi-encoder retrieval model, they asymmetrically encode queries and documents as separate dense vectors, and measure their pairwise relevance as the cosine similarity between the vectors. Owing to their strong in-domain performance and straightforward inference scheme, there has been a growing focus on improving their training. Studies demonstrate that large-scale unsupervised pair training utilizing in-batch negatives, followed by a small-scale triplet fine-tuning stage, significantly improves performance compared to a dense retriever trained solely on triplet data (li2023generaltextembeddingsmultistage; gunther2023jina). Other works have incorporated asymmetric task-specific instructions for queries and documents to further enhance performance (wang2024multilingual) and demonstrated the efficacy of using synthetically generated training data, including using diverse task instructions and machine translations, to further improve model representations. (wang2023improving; lee_gecko_2024)
@@ 2.2 Multi-Vector Retrieval
Multi-vector retrievers like ColBERT also employ a bi-encoder structure, but queries and passages are represented by a collection of smaller token embeddings rather than one large vector. As such, ColBERTv2’s training uses many of the same techniques as state-of-the-art single-vector models: cross-encoder distillation, multiple negatives per query, and self-mined hard negatives. Recent models have continued to improve on this training recipe, particularly for multilingual or non-English training. BGE-M3 (chen_bge_2024) adopts the two-stage pairs-to-triplets training pipeline, and does self-knowledge distillation, treating the combination of its sparse, dense, and multi-vector scores as the teacher score.
@@ 2.3 Multilingual Retrieval
Owing to the quality of English-based pre-trained models (BERT) and annotated data (MSMARCO), many advances in neural retrieval have been in a monolingual English setting. karpukhin_dense_2020; xiong_approximate_2020; khattab_colbert_2020 Researchers, however, have also made advances in non-English capabilities. On the modeling front, multilingual PLMs like mBERT (devlin_bert_2019) and later XLM-RoBERTa (conneau_unsupervised_2020) have expanded pre-training to include text in up to 100 languages, including in cross-language contexts. For multilingual retrieval, there are two approaches: natural and translated. Datasets like Mr-Tydi and MIRACL (mrtydi; zhang2023miracl) are built from human-generated and annotated queries, whereas mMARCO (bonifacio_mmarco_2022) is a collection of machine-translated copies of MSMARCO which inherit their judgments from the original dataset. The former method tends to be of higher quality and lacks the subtle distributional/idiomatic errors, dubbed "translationese", that the latter sometimes exhibits, however it costs more per example.
Recent multi-vector work has also proposed further modifications along the dimensions of architecture and data. ColBERT-XM louis_colbert-xm_2024 addresses the so-called curse of multilinguality (conneau_unsupervised_2020), the performance degradation of models pre-trained on too many tasks, with shared- and per-language parameters that allow for more robust zero-shot language transfer and post-hoc language extension. On the data approach, ColBERT-X (ecir2022colbert-x; lawrie_neural_2023; yang_distillation_2024) uses language-mixed batches of machine-translated English data, and BGE-M3 (chen_bge_2024) curates unsupervised and high-quality supervised corpora of diverse multilingual training data.
3 Evaluation Methodology
We evaluate Jina-ColBERT-v2 on three widely used benchmarks. For general English performance, we use the same subset of 14 retrieval and text-similarity tasks from the BEIR benchmark as in santhanam_colbertv2_2022. Additionally, we assess performance on the LoTTE benchmark, which focuses on long-tail queries, and the MIRACL and mMARCO benchmarks zhang2023miracl; bonifacio_mmarco_2022, which assess non-English retrieval performance. We report nDCG@10 for the BEIR and MIRACL collections, MRR@10 for mMARCO, and Success@5 for LoTTE. Scores are reported on the test split for BEIR, development split for MIRACL and mMARCO, and search test split for LoTTE. We use the same max query/document lengths as reported in santhanam_colbertv2_2022, and use the default (32/300) for MIRACL and mMARCO.
Due to the compute and time costs of indexing corpora containing tens of millions of documents, evaluating every model checkpoint on every task is not feasible. Therefore, we follow recent works (clavie_jacolbertv25_2024; merrick2024arcticembedscalableefficientaccurate) by comparing models’ quality on smaller sampled-corpus versions of HotpotQA, NQ, MS MARCO, and MIRACL (Chinese, French, German, Japanese, Spanish). These sampled corpora are constructed by combining the top 250 BM25-retrieved [^2] We use the standard pre-built Lucene indices in Pyserini (Lin_etal_SIGIR2021_Pyserini) for MIRACL found at https://github.com/castorini/pyserini/blob/master/docs/prebuilt-indexes.md, and use BM25s (bm25s) for BEIR. passages with all judged passages. We observe good agreement between the sampled-corpus evaluation scores and the full-fidelity ones when used to make binary model comparisons, but we leave a more rigorous analysis of this observation to future work. For the final model, we evaluate on the full version of every dataset.
@@ 4 Training Overview
    Triplet Training: Our model is further fine-tuned using retrieval examples in many languages with both positives and hard negatives, supervised by a highly-capable multilingual cross-encoder.
@@ 5.1 Backbone Improvements
Following many prior single- and multi-vector multilingual training efforts, we adopted XLM-RoBERTa as our backbone model due to its strong performance across various downstream tasks ecir2022colbert-x; louis_colbert-xm_2024; chen_bge_2024. We also enhance the XLM-RoBERTa architecture with flash attention (dao2023flashattention2) for efficiency and to replace its absolute positional embeddings with rotary positional embeddings (RoPE, su2023roformerenhancedtransformerrotary). To warm up its new positional embeddings, we continued pre-training the modified backbone with the same masked language modeling objective for 160,000 steps on the RefinedWeb dataset penedo2023refinedweb, a modern, high-quality corpus. During this pre-training phase, we set the maximum sequence length to 8,192 tokens with a rotary base of 10,000 and employed whole-word-masking devlin_bert_2019, masking out 30% of the tokens. We call this modified language model Jina-XLM-RoBERTa.
@@
Table 1: nDCG@10 Scores on BEIR datasets, grouped by task type (retrieval, question answering, and semantic text similarity) when using natural language instructions versus query/document marker tokens. Datasets marked with a * use the BM25-sampled corpus technique discussed in Section 3.
5.2 Multiple Linear Heads
To reduce index sizes, ColBERT includes a linear head that projects its token embeddings from the hidden dimension of its language model down to a lower dimension ($768\to 128$). Notably, BGE-M3’s multi-vector retrieval does not take this step, keeping its token embeddings at a full 1024 dimensions. To further reduce index size, we jointly train six linear heads with dimensions $d\in\{64,96,128,256,512,768\}$ using Matryoshka Representation Loss (MRL, kusupati_2022_matryoshka). This allows users to choose greater or lesser space efficiency, with an associated performance trade-off. Figure 1 shows the strong performance preservation of using reduced-dimension linear heads. Halving the token dimension ($128\to 64$) only causes its nDCG@10 to drop by 0.01 (1.59%). We unfortunately find that MRL’s weight-tying efficient variant (MRL-E), where losses are computed on truncations of the same vector does not preserve performance well.
@@ 6 Pair Training
To leverage an abundance of text pairs with rich semantic structures, we draw inspiration from common practices in single-vector embedding model training and begin by training on these text pairs, focusing on optimizing the embedding model’s performance on general semantic similarity tasks. In contrast, ColBERT typically trains directly on 32-way or 64-way triplets, consisting of a query, a positive passage, and multiple mined negatives.
@@ 6.1 Data Composition
Our pair training data consists of a broad range of weakly supervised datasets harvested from the web. We adjusted sampling rates across different languages and domains based on intuition, resulting in a set of 450 million weakly supervised, semantically related sentence pairs, question-answer pairs, and query-document pairs. Of these 450 million pairs, 50% are in English. The non-English language distribution is illustrated in Figure 2 in the Appendix. Our non-English pair-wise datasets contain a diverse collection of 29 major languages, including 3.0% code data, with 4.3% representing cross-lingual data.
@@ 6.2 Contrastive Loss
We utilize the same single-vector pair-training loss function as described in gunther2023jina.
Due to the typically symmetric nature of similarity measures, the loss is calculated in both directions. During the pair training stage, we set the temperature $\tau=0.02$ and used a peak learning rate of $5\times 10^{-5}$ with a warm-up period of 1,000 steps. The model was trained using the Adam optimizer for 100,000 steps with a global batch size of 16,384.
6.3 Task Instructions
Inspired by the use of instruction prefixes in single-vector works like su2022one, we experimented with adding task-specific natural language instructions for retrieval (RET), and question answering (QA), and semantic text similarity (STS). However, results in Table 1 show a generally negative effect across most BEIR datasets, with an average relative drop of -1.8% nDCG@10. We hypothesize that this is because instructions are not well-suited for late interaction models, which operate at the token level. Any embedding conditioning that the instructions might provide likely becomes less effective when aggregated at the token similarity level. Furthermore, these instructions occupy valuable space within the system’s fixed token capacity.
@@ 7.1 Data Composition
Our triplet dataset consists of 1) high-quality, human-annotated research datasets such as MSMARCO, DuReader, and MIRACL (bajaj2016msmarco; he2018dureaderchinesemachinereading; zhang2023miracl) with seven mined hard negatives per query and 2) high-quality datasets like MSMARCO and NQ translated from English into Chinese, French, German, Japanese, Russian and Spanish, following our previous work mohr2024multi and 3) synthetically generated datasets to address common failure modes of dense vector models such as negation and to cover niche domains like legal IR.
@@
The triplet dataset covers 14 widely used languages, with a strong emphasis on Arabic, Chinese, English, French, German, Japanese, Russian, and Spanish. We sample the datasets to create a language distribution similar to that used in pair training. In this distribution, outlined in Figure 3 in the Appendix, English accounts for 45.9% of the triplets, with 52.1% roughly evenly split between the mentioned high-resource non-English languages and a small 2% share for lower-resource languages.

### v2

# Jina-ColBERT-v2: A General-Purpose Multilingual Late Interaction Retriever

@@
Abstract
Multi-vector dense models, such as ColBERT, have proven highly effective in information retrieval. ColBERT’s late interaction scoring approximates the joint query-document attention seen in cross-encoders while maintaining inference efficiency closer to traditional dense retrieval models, thanks to its bi-encoder architecture and recent optimizations in indexing and search. In this paper, we introduce a novel architecture and a training framework to support long context window and multilingual retrieval. Leveraging Matryoshka Representation Loss, we further demonstrate that the reducing the embedding dimensionality from 128 to 64 has insignificant impact on the model’s retrieval performance and cut storage requirements by up to 50%. Our new model, Jina-ColBERT-v2, demonstrates strong performance across a range of English and multilingual retrieval tasks,
@@ 1 Introduction
Neural retrieval has gained popularity in recent years following the arrival of capable pre-trained language models (PLMs) (Devlin et al., 2019; Liu et al., 2019; Clark et al., 2020). Two types of approaches have been employed to apply PLMs in retrieval tasks. Sparse neural retrieval systems, such as SPLADE (Formal et al., 2021), represent texts as weighted bags of words. In contrast, dense retrievers encode queries and documents as high-dimensional vectors, capturing relevance signals through spatial relationships extending beyond exact term matching.
Most dense retrievers encode a query or document as a single vector, commonly the result of mean-pooling or the [CLS]-embedding over the transformer’s final layer token embeddings. The relevance is calculated between the query vector and the document vector. In comparison, multi-vector retrievers like ColBERT (Khattab and Zaharia, 2020) generalize this embedding process to maintain a embedding for each token, computing relevance scores as a late interaction over token embeddings.
Storing embeddings for each token, ColBERT models preserve the information at the most fine-grained granularity. With the late interaction functioned at the query time, ColBERT achieves better in-domain performance and tends to be more robust out-of-domain than single-vector similarity (Khattab and Zaharia, 2020). However, the context length of ColBERT is limited to 512 tokens which restricted the information stored in a single chunk.
Another drawback is that ColBERT takes more time to calculate because it needs to combine token interactions into one final score based on the embeddings of each token. Although it works well with the existing vector similarity infrastructure that makes single-vector methods efficient, ColBERT requires more storage space for the embeddings. To make the ColBERT usable in practice, the output dimensionality is restricted to be much smaller than the single-vector models.
Moreover, ColBERTv2 (Santhanam et al., 2022) is trained only on English MSMARCO triplets (Bajaj et al., 2016) and has a monolingual BERT backbone and is incapable of multilingual retrieval. Some previous works have made to extend the model to multilingual retrieval tasks. Louis et al. (2024) does this by using parameter extensions for each additional language, and Lawrie et al. (2023) trains solely on machine-translated English MSMARCO data to get effective multilingual performance. Due to the lack of diversity in the training corpus, these approaches come with significant trade-offs in terms of model usability and training data diversity.
Using the high-quality training resources, BGE-M3 (Chen et al., 2024) supports multi-vector retrieval with support to more than 100 languages and 8192 token context length. However, aiming at performing dense retrieval, multi-vector retrieval, and sparse retrieval simultaneously, it only produces embeddings with dimensionality as high as 1024. With such high-dimensional embeddings for each token, BGE-M3 has a limited utility for the multi-vector retrieval in practice.
@@
In this work, we propose Jina-ColBERT-v2, which introduces an improved training recipe for ColBERT models with the following features:
Weakly-supervised pretraining with RoPE:
Equipped with rotary position embedding and trained with large-scale unlabeled text pairs from various corpus with weakly-supervised contrastive pretraining, our model is able to offer high-quality multi-vector embeddings for 8k-token-long text. A second-stage finetuning with labeled data using supervised distillation is used to further boost its performances.
General multilingual performance:
@@
We train with data from a variety of high- and low-resource languages using both labeled and unlabeled data, including machine-translated training data, and show that this improves even out-of-domain multilingual performance. The diversity of the training data provide good performance in general on the multilingual retrieval tasks.
Inference-agnostic efficiency:
@@
We introduce multiple sizes of linear projection heads, jointly trained using the non-weight tying variant of Matryoshka Representation Loss Kusupati et al. (2022), enabling the selection of token embedding size at inference time with minimal performance degradation. We demonstrate that reducing the embedding dimensionality from 128 to 64 only introduces insignificant performance drop. Additionally, our flash-attention optimized backbone, Jina-XLM-RoBERTa  provides further performance boosts during inference.
@@ 2.1 Single-Vector Retrieval
Single-vector encoder models have demonstrated their potential as general-purpose embedding models across a number of downstream tasks (Muennighoff et al., 2023). When used in a bi-encoder retrieval model, they asymmetrically encode queries and documents as separate dense vectors, and measure their pairwise relevance as the cosine similarity between the vectors. Owing to their strong in-domain performance and straightforward inference scheme, there has been a growing focus on improving their training. Studies demonstrate that large-scale unsupervised pair training utilizing in-batch negatives, followed by a small-scale triplet finetuning stage, significantly improves performance compared to a dense retriever trained solely on triplet data (Li et al., 2023; Günther et al., 2023). Other works have incorporated asymmetric task-specific instructions for queries and documents to further enhance performance (Wang et al., 2024) and demonstrated the efficacy of using synthetically generated training data, including using diverse task instructions and machine translations, to further improve model representations. (Wang et al., 2023; Lee et al., 2024)
@@ 2.2 Multi-Vector Retrieval
Multi-vector retrievers like ColBERT also employ a bi-encoder structure, but queries and passages are represented by a collection of smaller token embeddings rather than one large vector. As such, ColBERTv2’s training uses many of the same techniques as state-of-the-art single-vector models: cross-encoder distillation, multiple negatives per query, and self-mined hard negatives. Recent models have continued to improve on this training recipe, particularly for multilingual or non-English training. BGE-M3 (Chen et al., 2024) adopts the two-stage pairs-to-triplets training pipeline, and does self-knowledge distillation, treating the combination of its sparse, dense, and multi-vector scores as the teacher score.
@@ 2.3 Multilingual Retrieval
Owing to the quality of English-based pre-trained models (BERT) and annotated data (MSMARCO), many advances in neural retrieval have been in a monolingual English setting Karpukhin et al. (2020); Xiong et al. (2020); Khattab and Zaharia (2020). Researchers, however, have also made advances in non-English capabilities. On the modeling front, multilingual PLMs like mBERT (Devlin et al., 2019) and later XLM-RoBERTa (Conneau et al., 2020) have expanded pre-training to include text in up to 100 languages, including in cross-language contexts. For multilingual retrieval, there are two approaches: natural and translated. Datasets like Mr-Tydi and MIRACL (Zhang et al., 2021, 2023b) are built from human-generated and annotated queries, whereas mMARCO (Bonifacio et al., 2022) is a collection of machine-translated copies of MSMARCO which inherit their judgments from the original dataset. The former method tends to be of higher quality and lacks the subtle distributional/idiomatic errors, dubbed "translationese", that the latter sometimes exhibits, however it costs more per example.
Recent multi-vector work has also proposed further modifications along the dimensions of architecture and data. ColBERT-XM Louis et al. (2024) addresses the so-called curse of multilinguality (Conneau et al., 2020), the performance degradation of models pre-trained on too many tasks, with shared- and per-language parameters that allow for more robust zero-shot language transfer and post-hoc language extension. On the data approach, ColBERT-X (Nair et al., 2022; Lawrie et al., 2023; Yang et al., 2024) uses language-mixed batches of machine-translated English data, and BGE-M3 (Chen et al., 2024) curates unsupervised and high-quality supervised corpora of diverse multilingual training data.
@@ 3 Training Overview
    Triplet Training: Our model is further finetuned using retrieval examples in many languages with both positives and hard negatives, supervised by a highly-capable multilingual cross-encoder.
@@ 4.1 Backbone Improvements
Following many prior single- and multi-vector multilingual training efforts, we adopted XLM-RoBERTa as our backbone model due to its strong performance across various downstream tasks Nair et al. (2022); Louis et al. (2024); Chen et al. (2024). To improve the efficiency, we enhance the XLM-RoBERTa architecture with flash attention (Dao, 2024).
In order to support long context window, We replace the absolute positional embeddings with rotary positional embeddings (RoPE, Su et al. (2023)). To warm up its new positional embeddings, we continued pre-training the modified backbone with the same masked language modeling objective for 160,000 steps on the RefinedWeb dataset Penedo et al. (2023), a modern, high-quality corpus. During this pre-training phase, we set the maximum sequence length to 8,192 tokens with a rotary base of 10,000 and employed whole-word-masking Devlin et al. (2019), masking out 30% of the tokens. We call this modified language model Jina-XLM-RoBERTa.
4.2 Multiple Linear Heads
To reduce index sizes, ColBERT includes a linear head that projects its token embeddings from the hidden dimension of its language model down to a lower dimension ($768\to 128$). Notably, BGE-M3’s multi-vector retrieval does not take this step, keeping its token embeddings at a full 1024 dimensions.
To further reduce index size, we jointly train six linear heads with dimensions $d\in\{64,96,128,256,512,768\}$ using Matryoshka Representation Loss (MRL, Kusupati et al. (2022)). This allows users to choose greater or lesser space efficiency, with an associated performance trade-off. Figure 1 shows the strong performance preservation of using reduced-dimension linear heads. Halving the token dimension ($128\to 64$) only causes its nDCG@10 to drop by 0.01 (1.59%). We unfortunately find that MRL’s weight-tying efficient variant (MRL-E), where losses are computed on truncations of the same token vector does not preserve performance well.
@@ 5 Pair Training
To leverage an abundance of text pairs with varying richness of semantic structure, we draw inspiration from common practices in single-vector embedding model training and begin with training on these text pairs, focusing on optimizing the embedding model’s performance on general semantic similarity tasks. This weakly-supervised stage is in contrast to previous ColBERT works, which typically starts with the triplet training on 32-way or 64-way retrieval triplets, consisting of a query, a positive passage, and multiple mined negatives.
@@ 5.1 Data Composition
Our pair training data consists of a broad range of weakly supervised datasets harvested from the web. We adjusted sampling rates across different languages and domains based on intuition, resulting in a set of 450 million weakly supervised, semantically related sentence pairs, question-answer pairs, and query-document pairs. Of these 450 million pairs, 50.0% are in English. Our non-English pair-wise datasets contain a diverse collection of 29 major languages, including 3.0% code data, with 4.3% representing cross-lingual data.
@@ 5.2 Contrastive Loss
We utilize the same single-vector pair-training loss function as described in Günther et al. (2023). Due to the often symmetric nature of our text pairs, the loss is calculated in both directions. During the pair training stage, we set the temperature $\tau=0.02$ and used a peak learning rate of $5\times 10^{-5}$ with a warm-up period of 1,000 steps. The model was trained using the Adam optimizer for 100,000 steps with a global batch size of 16,384.
@@ 6.1 Data Composition
Our triplet dataset consists of 1) high-quality, human-annotated research datasets such as MSMARCO, DuReader, and MIRACL (Bajaj et al., 2016; He et al., 2018; Zhang et al., 2023b) with seven mined hard negatives per query and 2) high-quality datasets like MSMARCO and NQ translated from English into Chinese, French, German, Japanese, Russian and Spanish, following our previous work Mohr et al. (2024) and 3) synthetically generated datasets to address common failure modes of dense vector models such as negation and to cover niche domains like legal IR.
@@
The triplet dataset covers 14 widely used languages, with a strong emphasis on Arabic, Chinese, English, French, German, Japanese, Russian, and Spanish. We sample the datasets to create a language distribution similar to that used in pair training. English accounts for 45.9% of the triplets, with 52.1% roughly evenly split between the mentioned high-resource non-English languages and a small 2.0% share for lower-resource languages.
6.2 Supervision Loss
Following ColBERTv2, we finetune our pair-trained checkpoint on samples with hard negatives using a KL divergence loss function to distill soft labels from the teacher model. For the teacher model, we use jina-reranker-v2-base-multilingual [^1] https://huggingface.co/jinaai/jina-reranker-v2-base-multilingual, a highly capable multilingual cross encoder.
7 Results
BEIR avg nf fi tc ar qu sd sf to db fe cf hp nq
BM25 44.0 32.5 23.6 65.6 31.5 78.9 15.8 66.5 36.7 31.3 75.3 21.3 60.3 32.9
ColBERTv2 49.6 33.7 35.4 72.6 46.5 85.5 15.4 68.9 26.0 45.2 78.5 17.6 67.5 52.4
answerai-v1 55.7 37.3 41.2 84.6 50.1 87.7 18.4 74.8 25.7 45.6 91.0 33.1 76.1 59.1
Ours 53.1 34.6 40.8 83.4 36.6 88.7 18.6 67.8 27.4 47.1 80.5 23.9 76.6 64.0
Table 1: Comparison of nDCG@10 scores between BM25, ColBERTv2, answer-colbert-small and Jina-ColBERT-v1 and Jina-ColBERT-v2 on the BEIR test set. The short names are used to refer to datasets and resources: nf for NFCorpus, fi for FIQA (Fact In Question Answering), tc for TREC-COVID (Text Retrieval Conference COVID), ar for Arguana, qu for Quora, sd for SciDocs, sf for SciFact, to for Webis-Touche, db for DBpedia-Entity, fe for FEVER (Fact Extraction and Verification), cf for Climate-FEVER, hp for HotpotQA, and nq for Natural Questions
We evaluate Jina-ColBERT-v2 on four widely used benchmarks, BEIR, LoTTE, and MIRACL and mMARCO. For general English performance, we use the same subset of 14 retrieval and text-similarity tasks from the BEIR benchmark as in Santhanam et al. (2022). Additionally, we assess performance on the LoTTE benchmark, which focuses on long-tail queries, and the MIRACL and mMARCO benchmarks Zhang et al. (2023b); Bonifacio et al. (2022), which assess non-English retrieval performance. We report nDCG@10 for the BEIR and MIRACL collections, MRR@10 for mMARCO, and Success@5 for LoTTE. Scores are reported on the test split for BEIR, development split for MIRACL and mMARCO, and search test split for LoTTE. We use the same maximum query/document lengths as reported in Santhanam et al. (2022), and use the default (32/300) for MIRACL and mMARCO.
LoTTE avg Life. Rec. Wri. Sci. Tech.
BM25 67.8 80.2 68.5 74.7 53.6 61.9
ColBERTv2 72.0 84.7 72.3 80.1 56.7 66.1
Ours 76.4 87.0 77.6 83.8 60.5 73.0
Table 2: Comparison of Success@5 of various models across different LoTTE search query subsets.
Table 1 shows Jina-ColBERT-v2’s strong English performance compared to ColBERTv2, while still trailing the monolingual answerai-colbert-small-v1. Notably, however, we perform well below ColBERTv2 on ArguAna (ar), which we might attribute to either its unusual task: counterargument retrieval being at odds with our retrieval-heavy triplet training data distribution, or as an indication of the limitation of our stronger augmentation attention (discussed in Section 8.4) when applied to much longer (300 token) queries. Similarly for LoTTE, we see in Table 2 an improvement over ColBERTv2.
MIRACL avg ar bn de es en fa fi fr hi id ja ko ru sw te th yo zh
BM25 38.5 48.1 50.8 22.6 31.9 35.1 33.3 55.1 18.3 45.8 44.9 36.9 41.9 33.4 38.3 49.4 48.4 40.6 18.0
mDPR-ZS 41.8 49.9 44.3 49.0 47.8 39.4 48.0 47.2 43.5 38.3 27.2 43.9 41.9 40.7 29.9 35.6 35.8 39.6 51.2
mDPR-FT 62.7 72.5 68.4 - 48.8 56.5 59.3 71.4 58.9 51.6 49.6 64.2 59.0 59.7 68.5 80.4 69.5 - 65.0
Ours 62.3 75.3 75.0 50.4 53.8 57.0 56.3 74.0 54.1 60.0 54.7 63.2 67.1 64.3 49.9 74.2 77.2 62.3 52.3
Table 3: Comparison of nDCG@10 scores for BM25, mDPR-ZeroShot (ZS), mDPR-FineTuned (FT), and Jina-ColBERT-v2 models on the MIRACL dev set across various languages.
Table 3 compares Jina-ColBERT-v2 to BM25, mDPR, and BGE-M3. While we handily outperform BM25 and zero-shot mDPR (Zhang et al., 2023b) as expected, our model is slightly outperformed by the finetuned mDPR (Zhang et al., 2023a). For context, each mDPR-FT is only tuned on one language, rather than many like ours which may suffer to some extent from the curse of multilinguality.
mMARCO avg ar de nl es fr hi id it ja pt ru vi zh
BM-25 13.9 11.1 13.6 14.0 15.8 15.5 13.4 14.9 15.3 14.1 15.2 12.4 13.6 11.6
ColBERT-XM 25.4 19.5 27.0 27.5 28.5 26.9 23.8 26.3 26.5 24.1 27.6 25.1 22.6 24.6
Ours 31.3 27.2 33.1 33.0 34.1 33.5 30.9 31.9 33.7 27.6 33.7 29.8 28.7 30.2
Table 4: Comparison of mRR@10 scores between BM25, ColBERT-XM and Jina-ColBERT-v2 models on the mMARCO dev set across various languages.
8 Ablation Studies
In this section we present short ablation studies on modifications to three various aspects of ColBERT modeling and training.
8.1 Efficient Evaluation
Due to the compute and time costs of indexing corpora containing tens of millions of documents, evaluating every model checkpoint and ablation on every task is not feasible. Therefore, we follow recent works (Clavié, 2024; Merrick et al., 2024) by comparing models’ quality on smaller sampled-corpus versions of HotpotQA, NQ, MS MARCO, and MIRACL (Chinese, French, German, Japanese, Spanish). These sampled corpora are constructed by combining the top 250 BM25-retrieved [^2] We use the standard pre-built Lucene indices in Pyserini (Lin et al., 2021) for MIRACL found at https://github.com/castorini/pyserini, and use BM25s (Lù, 2024) for BEIR. passages with all judged passages. We observe good agreement between the sampled-corpus evaluation scores and the full-fidelity ones when used to make binary or ranking-based model comparisons, but we leave a more rigorous analysis of this observation to future work. We only use the sampled corpora for ablation studies. For the final model, we evaluate on the full version of every dataset.
8.2 Task Instructions
Inspired by the use of instruction prefixes in single-vector works like Su et al. (2022), we experimented with adding task-specific natural language instructions for retrieval (RET), and question answering (QA), and semantic text similarity (STS). However, results in Table 5 show a generally negative effect across most BEIR datasets. We hypothesize that this is because instructions are not well-suited for late interaction models, which operate at the token level. Any embedding conditioning that the instructions might provide likely becomes less effective when aggregated at the token similarity level. Furthermore, these instructions occupy valuable space within the system’s fixed token capacity.
RET QA STS
nf tc sf to db fe cf ms* fq hp* nq* ar qu sd
Mark. 32.4 59.3 67.9 19.3 35.3 67.1 18.3 34.4 37.5 25.9 40.8 37.5 86.1 18.4
Inst. 32.9 63.2 67.5 18.8 33.9 64.4 16.7 34.0 37.1 24.9 42.9 34.2 86.0 17.9
Table 5: nDCG@10 scores on BEIR datasets, grouped by task type (retrieval, question answering, and semantic text similarity) when using natural language instructions versus query/document marker tokens. Datasets marked with a * use the BM25-sampled corpus technique discussed in Section 8.1.
8.3 Score Normalization
BEIR MIRACL
tc hp nq ms de es fr ja zh
with score normalization 80.1 36.4 56.6 45.1 57.7 39.3 51.3 61.8 62.5
without score normalization 78.7 36.6 58.0 45.4 57.3 40.6 50.7 63.4 63.2
Table 6: nDCG@10 scores for score normalization experiments on a retrieval-oriented subset of BEIR and MIRACL tasks. Results are performed on the BM25-sampled versions of all datasets presented except TREC-COVID (tc).
Recently, Clavié (2024) applied min-max normalization to both the student and teacher scores before computing the KL loss. This adjustment brings the score distributions of the ColBERT model and its CE teacher into closer alignment, as the original score distribution for ColBERT theoretically ranges from zero to the number of query tokens, and is model-dependent for the teacher CE. Our experiment presented in Table 6, however, shows this method to have inconclusive benefit to nDCG@10 on the BEIR and MIRACL datasets when applied to our model. We consider this result to be understandable given Clavié (2024)’s very small observed effect.